# AKRMap: Adaptive Kernel Regression for Trustworthy Visualization of Cross-Modal Embeddings

### v1

@@ Appendix A Performance in Traditional Neighborhood Preservation Objective
Traditional dimensionality reduction methods often seek to preserve the neighborhood in high-dimensional space in the visualization scatterplot. Since the target of our method is not to optimize the neighborhood preservation, we do not expect our method to be better than traditional dimensionality reduction method in this regard. In Table 2, we provide statistics of traditional neighborhood trustworthiness metric (Kaski et al., 2003) to show that our method does not significantly harm the neighborhood property. Furthermore, we note that in our application where we focus on continuous metric, cluster separation is not a meaningful target since we are not looking at discrete categories as in classification tasks. As shown in Figure 8, for the HPSv2 metric on HPD dataset, even in scatterplot view AKRMap can reveal the distribution of metric significantly better than other traditional DR methods.
@@
Figure 8: Comparison of scatterplots produced by different DR methods.
@@ Appendix C RBF Kernel Settings for Traditional DR Methods
Table 6: Impact of manually setting smaller bandwidth.
@@ Appendix D Zoom in by different grid numbers
This section shows the zoom in effect of AKRMap when increasing the grid number. As shown in Figure 10, AKRMap can accurately estimate the local distribution of cross-modal metric and adds detail on demand. Specifically, with a small grid number of 30 by 30, AKRMap can already achieve the level of detail close to most baseline methods with grid number of 500 by 500. When increasing the grid number, we can see in Figure 10 that our method can accurately show the detail contour of metric distribution.
@@ Appendix E Hyperparameter
In this section we discuss some hyperparameter trade-off in our method with additional experimental results.
@@ Appendix F Test on other modality embeddings
We further perform quantitative experiments on CLIP text-video embeddings on MSR-VTT dataset which consists of 200,000 text-video pairs, as well as text-audio embeddings and image-audio embeddings on Flickr 8k Audio Caption Corpus dataset consisting of 40,000 pairs.
@@
Table 10: CLIP text-video embeddings mapping accuracy on MSR-VTT datasets
@@
Appendix G Weighting balancing mechanism
@@
$L_{2}=w(MSE_{r},\mu_{1})\cdot\lambda MSE_{r}+\cdot KL,$ (12)
@@ Appendix H Interactive features evaluation
We conduct a user study and record the response time of our interactive features including zoom and overlay.

### v2

@@ Appendix A Performance in Traditional Neighborhood Preservation Objective
Traditional dimensionality reduction methods often seek to preserve the neighborhood in high-dimensional space in the visualization scatterplot. Since the target of our method is not to optimize the neighborhood preservation, we do not expect our method to be better than traditional dimensionality reduction method in this regard. In Table 2, we provide statistics of traditional neighborhood trustworthiness metric (Kaski et al., 2003) to show that our method does not significantly harm the neighborhood property. Furthermore, we note that in our application where we focus on continuous metric, cluster separation is not a meaningful target since we are not looking at discrete categories as in classification tasks. As shown in Figure 8, for the HPSv2 metric on HPD dataset, even in scatterplot view AKRMap can reveal the distribution of metric significantly better than other traditional DR methods. Here we also note that we deliberately select larger perplexity for the t-SNE method because of the large scale of the HPD dataset, and we show in Figure 9 that smaller perplexity will lead to worse results.
@@
Figure 8: Comparison of scatterplots produced by different DR methods for HPSv2.
Figure 9: Results of smaller perplexity for t-SNE on HPSv2.
@@ Appendix C RBF Kernel Settings for Traditional DR Methods
Table 6: Impact of manually setting smaller bandwidth for Figure 10.
@@ Appendix D Zoom in by Different Grid Numbers
This section shows the zoom-in effect of AKRMap when increasing the grid number. As shown in Figure 11, AKRMap can accurately estimate the local distribution of cross-modal metric and adds detail on demand. Specifically, with a small grid number of 30 by 30, AKRMap can already achieve the level of detail close to most baseline methods with grid number of 500 by 500. When increasing the grid number, we can see in Figure 11 that our method can accurately show the detail contour of metric distribution.
@@ Appendix E Hyperparameter
In this section we discuss some hyperparameter trade-off in our method with additional experimental results. First, we show the trade-off effect by different $\lambda$ value to weight the regression loss and KL loss, with test (out-of-sample) quantitative results shown in Table 7, and visualization effects shown in Figure 12. Next, we show the scatterplot view of our ablation studies in Figure 13 to demonstrate the effect of our proposed component in large-scale DR point plot. Finally, in Figure 14, we illustrate some qualitative and quantitative effect of setting the train-val weights $w_{1}$ or $w_{2}$ to be zero to demonstrate the necessity of the train-val balance scheme in our kernel regression guidance loss.
@@ Appendix F Test on Other Modality Embeddings
We further perform quantitative experiments on CLIP text-video embeddings on MSR-VTT dataset (Xu et al., 2016) which consists of 200,000 text-video pairs (Table 10), as well as text-audio embeddings and image-audio embeddings on Flickr 8k Audio Caption Corpus dataset (Harwath & Glass, 2015) consisting of 40,000 pairs (Table 9 and Table 8). Overall, the results indicate that our method can improve upon traditional DR in test set mapping accuracy.
@@
Table 10: CLIP text-video embeddings mapping accuracy on MSR-VTT datasets.
@@
Appendix G Weight Balancing Mechanism
@@
$L_{2}=w(MSE_{r},\mu_{1})\cdot\lambda MSE_{r}+\cdot KL.$ (12)
The experimental results for different $\mu$ and $\mu_{1}$ settings are presented in Table 11.
@@ Appendix H Interactive Features Evaluation
We conduct a user study among 12 users on our interactive features including zoom and overlay. They complete a 7 Likert-scale survey after using our interactive visualization of HPSv2 score on HPD dataset. As shown in Figure 15, users generally appreciate our interactive features in several usability aspects including effectiveness, ease of use, interpretability, and future use.
@@
Appendix I Convergence Properties of the Adaptive Generalized Kernel Regression
In this appendix, we summarize the theoretical conditions and practical considerations related to the convergence of our adaptive generalized kernel regression method, which extends the classical Nadaraya-Watson kernel regression framework.
To guarantee convergence in terms of mean squared error (MSE), the following conditions are essential:
  • 1) Integrability of the generalized kernel on $\mathbb{R}^{2}$: The kernel is defined as:
    $K(x)=(1+\alpha|x|^{2\beta})^{-1}.$ (13)
    For $K(x)$ to be integrable and finite over $\mathbb{R}^{2}$, parameters must satisfy $\alpha>0$ and $\beta>1$, ensuring sufficient decay of the kernel function at infinity.
  • 2) Growth behavior of the parameter $\alpha$ with sample size $N$: The parameter $\alpha$ must effectively increase as the sample size $N$ grows, analogous to the classical kernel bandwidth $h$ decreasing with larger $N$. This condition is critical for establishing consistency and uniform convergence of the estimator.
  • 3) Smoothness of the regression target function $s(x)$: The target function $s(x)$ should satisfy smoothness assumptions such as Lipschitz continuity. This is a standard requirement in kernel regression theory to control the bias term and facilitate uniform convergence as bandwidth parameters are tuned.
In our training procedure, we enforce the non-negativity of parameters $\alpha$ and $\beta$ through reparameterization (by squaring them) and include them within a set of learnable parameters updated via backpropagation. Intuitively, this process pushes the network to automatically identify suitable decay rates for the kernel in each epoch or mini-batch, adapting progressively to match the magnitude scale of the Nadaraya-Watson kernel. Empirical experiments demonstrate that the learned parameter values ($\alpha$, $\beta$) converge successfully. For example, the values for HPSv2 are (68.57, 1.61), for PickScore are (59.13, 1.35), for CLIPScore are (104.70, 3.18), and for Aesthetic Score are (74.95, 1.11). These results indicate that our data-driven approach effectively satisfies conditions (1) and (2) mentioned above.
In practical multimodal scenarios, although the target function might not strictly satisfy Lipschitz continuity, numerical experiments indicate that kernel regression methods remain robust as long as the target distribution does not exhibit extreme irregularities, such as highly discontinuous or severely jagged patterns. In other words, even piecewise smooth or locally continuous target functions can yield stable and accurate kernel regression estimates in practice. Specifically, in our scenario involving high-dimensional to low-dimensional mappings, there exists an continuous relationship between evaluation metrics and the projected cross-modal embeddings. For example, CLIPScore, which is computed using cosine similarity between image and text embeddings, is an obvious continuous function between the original high-dimensional embeddings and the metric. Furthermore, our parametric projection network is an explicit continuous and differentiable function. The Implicit Function Theorem can then ensure the continuity of the metric $w.r.t$ the projected 2D embeddings. Thus, our practical setup implicitly fulfills the smoothness assumption (condition 3) required for convergence.
Furthermore, the mapping function $s(x)$ is continuously updated during dimensionality reduction (projection). Our adaptively updated generalized kernel not only naturally accommodates complex distributions encountered in practice but also ensures consistency in metric spaces through supervised kernel regression. Regarding projection stability, our kernel regression loss can be regarded as a regularization term that enforces the global structure of the projection, as it explicitly pushes the projected sample points to regress a global metric distribution.